# SiLVR: Scalable Lidar-Visual Radiance Field Reconstruction with Uncertainty Quantification

### v1

@@ Abstract
[…]
Index Terms:
Mapping, Sensor Fusion, RGB-D Perception, Neural Radiance Field (NeRF), Uncertainty Estimation
@@
This project has been partly funded by the Horizon Europe project Digiforest (101070405). Maurice Fallon is supported by a Royal Society University Research Fellowship. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Accepted Manuscript version arising.
@@ I Introduction
Dense 3D reconstruction is a core component of a range of robotics applications such as industrial inspection and autonomous navigation. Common sensors used for reconstruction include cameras and lidars. Camera-based reconstruction systems use techniques including Structure-from-Motion (SfM [1]) and Multi-View Stereo (MVS) [2] to produce dense textured reconstructions. However, these approaches rely on good lighting conditions and can require exhaustive data collection to capture data from diverse viewpoints. They also struggle with textureless areas such as bare walls, ceilings and floors. A lidar sensor provides accurate geometric information at long range—as it actively measures distances to surfaces. This makes lidar desirable for large-scale outdoor environments. However, lidar scans are much sparser than camera images. They also do not capture colour which is important for applications such as virtual reality and 3D asset generation.
@@
As with traditional vision-based reconstruction methods, NeRF struggles to estimate accurate geometry in locations where there is limited multi-view input (i.e. when images are only taken from a single direction) or sparse texture. These challenges affect not only volume-density-based fields, but also SDF-based fields [5] and 3D Gaussian representations [4], as can be seen later in Fig. 7. Autonomous systems commonly encounter these situations—for example, a robot reconstructing a wall with uniform colour when travelling directly towards it. In this paper, we show how this problem can be addressed by using lidar sensing to provide accurate geometric measurements of these types of textureless objects. In addition, the use of lidar data can reduce the need to capture varied camera views. It is impractical for an inspection robot (e.g., Spot quadruped on an industrial facility) to execute object-centric trajectories just to improve visual reconstruction. This motivates the development of a reconstruction system that fuses these sensors.
@@
The reconstruction task becomes even more challenging when the scene is large-scale (e.g., urban districts). Running an incremental SfM algorithm such as COLMAP [1] for a large-scale scene is time-consuming, and is not guaranteed to succeed in registering all input images (especially if parts of the scene have poor lighting), which then could lead to an unreconstructed region in the map. As the scene and the dataset size grow, the model capacity of a NeRF and memory constraints become a bottleneck. Simply increasing the size of the model parameters (e.g., hash table size in Instant-NGP [6] or the number of 3D Gaussians [4]) can exceed available computer memory when working on larger scenes. This motivates the development of a scene partitioning strategy. Some existing methods require manually partitioning using heuristics [7] or evenly partitioning the scene using a grid [8]. The limitation of a simple grid-based partition is that the view orientation and visibility are not considered, while it is important to consider these factors while carrying out clustering for visual reconstruction. For example, an image taken from a corridor outside of a room but looking into it ought to be considered part of a submap of that room.
The reconstructed 3D map from a NeRF pipeline can contain artefacts due to unconstrained views, and removing them can be challenging. The implicit representation used in NeRF, while providing tremendous model size compression compared to explicit representations such as 3DGS [4], can generate reconstruction artefacts outside of the observed area. The standard NeRF formulation has no notion of uncertainty for parts of the reconstruction that are unreliable. The regions which can be confidently reconstructed by a camera or a lidar are usually complementary: regions that have uniform texture or lack multi-view constraints often have high visual uncertainty, and regions that are outside of the lidar Field-of-View or beyond the lidar’s sensing range often have high lidar uncertainty. These reconstruction artefacts make it challenging when merging submaps, as each submap will contain noisier reconstruction on its boundary as the measurement density falls off. This can lead to artefacts in the overall merged map. These issues motivate the need for a principled approach to quantifying the reconstruction uncertainty of NeRF maps.
@@
Compared to prior NeRF-based reconstruction systems that incorporate lidar [12] or depth cameras [13, 11], our key innovation is a rigorous study of how to quantify uncertainty in the resultant reconstruction, which enables improved reconstruction accuracy as well as facilitating downstream tasks such as active mapping and navigation. After training, SiLVR computes the epistemic uncertainty of the NeRF with the Laplace approximation (LA) [14] and Fisher-Information-approximated Hessian matrix. As an efficient alternative to ensemble learning (which requires multiple iterations of training of the NeRF model and is time-consuming), we build upon the formulation of the perturbation field proposed in BayesRays [15] and use the spatial variance of the field as the measure of epistemic uncertainty. This is used to filter out reconstruction artefacts which improves the final reconstruction accuracy. We also adapt a previously developed lidar SLAM algorithm [16, 17] to bootstrap the SfM component accelerating its operation by up to 70%) and to enforce an accurate metric scale. For mapping large-scale building complexes or a city block, we adopt a submapping approach and apply graph partitioning algorithm [18] with visibility information to divide the complete trajectory into submaps. We study how the use of visibility information allows the submaps to be more self-contained and to have fewer artefacts at their boundaries compared to methods that only consider spatial information [12].
@@
  • Epistemic uncertainty quantification of the multi-modal NeRF pipeline which considers lidar and visual uncertainties differently. This allows us to filter parts of the reconstruction that are unreliable by directly using sensor-specific characteristics. To do this, we extended BayesRays [15]’s vision-only uncertainty framework to also support lidar depth. This allows us to identify areas with reliable reconstruction (e.g., where there are visual features or abundant lidar measurements) and unreliable areas (e.g., uniformly-textured surfaces which have also not been scanned by lidar). Filtering based on the uncertainty estimates improves the reconstruction accuracy. This is a more principled approach compared to the surface-density-based filtering [12].
@@ II-A Classical 3D Reconstruction
Lidars and cameras are the two main modalities used in robotic perception and specifically for 3D reconstruction. In this section, we review classical lidar-based and vision-based reconstruction methods. For each sensor modality, trajectory estimation is a typical first step in a reconstruction pipeline. Then, we review strategies for extending these methods to large-scale scenes.
@@
Lidar is the dominant sensor used for accurate 3D reconstruction of large-scale environments [20, 21]. It actively transmits laser pulses to measure ranges and as a result is accurate up to ranges of as much 100m. With these distance measurements, Lidar odometry typically uses a variant of Iterative Closest Point (ICP), and often integrates high-frequency IMU measurements [22, 23, 24, 16]. Small errors in odometry can accumulate over time resulting in trajectory drift. This drift can be mitigated when a sensor revisits a previous place and detects loop closure with pose graph optimisation [25] which allows a consistent map to be maintained. After registering all the lidar scans, the (surface) reconstruction problem is then to fuse discrete observations into a map. Example map representations include surfels [26, 27], voxels [28, 29, 30, 31] and wavelets [32]. Despite its advantages, lidar has its own limitations. Lidar is much more expensive than cameras, while the measurements are typically much sparser than camera images. The measurements have inherent noise with ranging errors in the order of centimetres, which makes it difficult to reconstruct small objects accurately in indoor scenes. Finally, lidar data contains no texture or colour, so the final reconstruction is only geometric and cannot be used for applications such as view synthesis, which requires texture.
@@
Alternatively, large-scale textured reconstructions can be recovered from camera images alone via SfM. Given the correspondences between images, a SfM system [1] can optimise a set of camera poses, camera intrinsics, and 3D sparse feature points. This can then be used by a MVS system [33] to compute dense depth for each frame and in turn to create a dense 3D point cloud. Compared to lidar, cameras are much more affordable, and also provide texture and colour. However, the performance of visual reconstruction method depends on environmental conditions, and the quality of feature matching. This makes the resultant reconstruction less reliable in scenes that contain repetitive patterns, low-texture surfaces, dynamic objects, poor lighting conditions and non-Lambertian materials.
@@
When the scene to be reconstructed is large-scale (e.g., urban districts or multi-room indoor environments), computer memory becomes a limiting factor. Attempting to map a large scene while constraining output map size will result in a lower resolution reconstruction, which results in a map that lacks detail. A common strategy to extend dense reconstructions to large-scale areas is to divide the scene into submaps [34]. For visual reconstruction with many thousands of images [35], a submapping approach can significantly reduce computation time and memory requirements. One approach used in large-scale MVS is the submap partitioning [36] which groups images into clusters while not degrading the final resultant reconstruction. After partitioning, each submap should be filtered and merged into one unified model. For online lidar mapping systems, the motivation for using submapping techniques is to accommodate loop closure corrections into the already-built map (Occupancy grid or TSDF). These systems construct submaps that are attached to a pose graph [37, 38, 39], and can deform each submap by reoptimising the pose graph upon loop closure.
@@ II-B Radiance Field Representation
Neural Radiance Fields (NeRF) was first proposed in the seminal paper from Mildenhall et al. [3]. The technique uses a multilayer perceptron (MLP) to represent a continuous radiance field, and uses differentiable volume rendering to reconstruct novel views. It minimises the photometric error between the rendered image and the input image, which implicitly achieves multi-view consistency. NeRF and its many variants use frequency encoding [40] to encode spatial coordinates, but these suffer from long training times, typically a few hours per scene. Alternative explicit representations of radiance fields, including dense voxel-grids with trainable per-vertex features [41, 6] and more recently 3D Gaussians [4] are shown to accelerate the training, at the cost of being more memory intensive. Octree or sparse-grid structures [42, 6] can reduce memory usage by pruning grid-features in empty space. Our work is built upon Nerfacto from the open-sourced Nerfstudio project [9]. It incorporates the main features from other NeRF works [6, 43, 44] that have been found to work well with real data.
@@
Neural field representations have been used for lidar-based mapping [49, 50, 51], showing promise in generating more complete and compact reconstructions than traditional methods. While these works also build upon implicit map representations, they do not use visual data to build the map. Our system uses visual information and multi-view geometry constraints. Because of this, it can reconstruct regions outside of the lidar’s field-of view.
@@ III-A Radiance Field Formulation
where $\mathbf{c}_{i}$ is the colour of the $i$-th point sample along the ray, and its weighting coefficient $w_{i}$ is defined by
@@
where $\delta_{i}$ is the distance between adjacent samples.
@@
$\mathcal{L}_{\text{Colour}}=\sum_{\mathbf{r}\in\mathcal{D}}||\hat{\mathbf{C}}(\mathbf{r})-\mathbf{C}(\mathbf{r})||^{2}$ (3)
@@
Our system extends upon Nerfacto, which is a specific pipeline implemented within the Nerfstudio framework [9]. Nerfacto’s rendering quality is comparable to state-of-the-art methods such as MipNeRF-360 [43] while achieving a substantial acceleration in reconstruction speed as it also incorporates efficient hash encoding which was proposed by the authors of Instant-NGP [6]. The scene contraction, proposed in [43], is also used to improve memory efficiency and to represent scenes with high-resolution content near the input camera locations. The contraction function non-linearly maps any point in space into a cube of side length 2, and represents the scene within this contracted space. In real-world outdoor environments, there is often large variation in exposure and lighting conditions. Because of this we use a per-frame appearance encoding for each image, similar to [45, 44].
@@ III-B Bayesian Interpretation of the NeRF training
The NeRF reconstruction i.e. the radiance field function $f$ is deterministic, and has no explicit notion of uncertainty. In practice, different parts of the NeRF reconstruction are inherently more uncertain or less reliable. For example, ill-conditioned visual constraints from non-textured areas or insufficient visual parallax can cause the visual reconstruction accuracy to deteriorate. Another example specifically relevant to our work is that a surface is more uncertain if it has only been sparsely scanned by the lidar with limited range and Field-of-View compared to a surface that is densely scanned. Quantifying these uncertainties associated with the NeRF reconstruction allows one to identify the unreliable reconstruction and filter them accordingly, hence improving reconstruction accuracy. It can also enable downstream tasks such as active mapping and navigation [57].

### v3

@@ Abstract
We present a neural radiance field (NeRF) based large-scale reconstruction system that fuses lidar and vision data to generate high-quality reconstructions that are geometrically accurate and capture photorealistic texture. Our system adopts the state-of-the-art NeRF representation to additionally incorporate lidar. Adding lidar data adds strong geometric constraints on the depth and surface normals, which is particularly useful when modelling uniform texture surfaces which contain ambiguous visual reconstruction cues. A key contribution of this work is a novel method to quantify the epistemic uncertainty of the lidar-visual NeRF reconstruction by estimating the spatial variance of each point location in the radiance field given the sensor observations from the cameras and lidar. This provides a principled approach to evaluate the contribution of each sensor modality to the final reconstruction. In this way, reconstructions that are uncertain (due to, e.g., uniform visual texture, limited observation viewpoints, or little lidar coverage) can be identified and removed. Our system is integrated with a real-time pose-graph lidar SLAM system which is used to bootstrap a Structure-from-Motion (SfM) reconstruction procedure. It also helps to properly constrain the overall metric scale which is essential for the lidar depth loss. The refined SLAM trajectory can then be divided into submaps using Spectral Clustering to group sets of co-visible images together. This submapping approach is more suitable for visual reconstruction than distance-based partitioning. Our uncertainty estimation is particularly effective when merging submaps, as their boundaries often contain artefacts due to limited observations. We demonstrate the reconstruction system using a multi-camera, lidar sensor suite in experiments involving both robot-mounted and handheld scanning. Our test datasets cover a total area of more than $\text{20,000}~\text{m}^{2}$, including multiple university buildings and an aerial survey of a multi-storey building. Quantitative evaluation is provided by comparing with maps produced by a commercial tripod scanner. The code and dataset are available at https://dynamic.robots.ox.ac.uk/projects/silvr/.
@@
This project has been partly funded by the Horizon Europe project Digiforest (101070405) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(No. RS-2024-00461409). Maurice Fallon is supported by a Royal Society University Research Fellowship. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Accepted Manuscript version arising.
@@ I Introduction
Dense 3D reconstruction is a core component of a range of robotics applications such as industrial inspection and autonomous navigation. Common sensors used for reconstruction include cameras and lidars. Camera-based reconstruction systems use techniques including Structure-from-Motion (SfM) [1] and Multi-View Stereo (MVS) [2] to produce dense textured reconstructions. However, these approaches rely on good lighting conditions and can require exhaustive data collection to capture data from diverse viewpoints. They also struggle with textureless areas such as bare walls, ceilings and floors. A lidar sensor provides accurate geometric information at long range—as it actively measures distances to surfaces. This makes lidar desirable for large-scale outdoor environments. However, lidar scans are much sparser than camera images. They also do not capture colour which is important for applications such as virtual reality and 3D asset generation.
@@
As with traditional vision-based reconstruction methods, NeRF struggles to estimate accurate geometry in locations where there is limited multi-view input (i.e., when images are only taken from a single direction) or sparse texture. Autonomous systems commonly encounter these situations—for example, an inspection robot moving only in a forward direction might only obtain visual observations from parallel viewpoints, which makes triangulation of visual features more difficult. Additionally, a robot operating in man-made environments often encounters objects such as textureless walls which are difficult to reconstruct using only vision. As a result, radiance field reconstruction of feature-less regions (e.g., the ground reconstruction using Nerfacto in Fig. 6; the sky and clouds reconstructed as floating points on the top of the scene [5]) and objects with limited viewpoints (e.g., reconstruction of a wall viewed from a single direction shown in Fig. 12) are often inaccurate. These challenges affect not only volume-density-based fields such as NeRFs, but also SDF-based radiance fields such as NeuS [6] and 3D Gaussian representations [4] (e.g., NeuSfacto reconstruction in Fig. 6 and elongated Gaussians described in MonoGS [7]). In addition, the implicit representation used in NeRF, while providing tremendous model size compression compared to explicit representations such as 3DGS [4], can generate reconstruction artefacts in unobserved regions of space. This is due to NeRF formulation as a continuous volumetric field over $\mathbb{R}^{3}$, which enables it to produce outputs even in areas lacking observation. Unlike classical visual SLAM/MVS methods (such as MonoSLAM [8] and OpenMVS) which estimate the uncertainties for the visual reconstructions to tackle reconstruction artefacts, the NeRF representation itself has no notion of uncertainties. These factors limit the use of radiance fields in real-world robotic applications where 3D reconstructions have to be reliable.
The challenges of 3D reconstruction using only visual sensing can be mitigated by fusing range sensing devices such as lidar which provides accurate geometric measurements regardless of object texture. Several recent reconstruction systems [9, 10, 11] have adopted a lidar-visual radiance field approach, since both RGB colour and depth measurements can be obtained (in a differentiable manner) from the radiance field and then supervised by lidar and camera data. However, lidar measurements are sparse, and have limited range and field-of-view (FOV). As a result, not all parts of the scene can be scanned by a lidar sensor. For example, the lidar in Fig. 10 has a limited FOV, and the top of the buildings are not scanned by it. As another example, the ground shown in the bottom left of Fig. 4 is beyond the lidar’s sensing range. For these regions, a lidar-visual radiance field reconstruction relies mainly on visual data, and the accuracy then depends on the conditioning of visual reconstruction as discussed in the previous paragraph. Because of this, it is crucial to properly quantify the individual contributions of visual and lidar information in the resultant lidar-visual radiance field reconstruction. However, this issue is relatively under-explored in the literature.
@@
The reconstruction task becomes even more challenging when the scene is large-scale (e.g., urban districts). It is time-consuming to estimate sensor trajectories by running an incremental SfM algorithm such as COLMAP [1] for a large-scale scene. In addition, SfM is not guaranteed to succeed in registering all input images (especially if parts of the scene have poor lighting), which could then lead to an unreconstructed region in the map. As the scene and the dataset size grow, the model capacity of a NeRF and memory constraints become a bottleneck. Simply increasing the size of the model parameters (e.g., hash table size in Instant-NGP [12] or the number of 3D Gaussians [4]) can exceed available computer memory when working on larger scenes. This motivates the development of a scene partitioning strategy. Some existing methods require manually partitioning using heuristics [13] or evenly partitioning the scene using a grid [14]. The limitation of a simple grid-based partition is that the view orientation and visibility are not considered. These factors are, however, important to consider when carrying out clustering for visual reconstruction. For example, an image taken from a corridor outside of a room but looking into it ought to be considered part of a submap of that room.
@@
Compared to prior NeRF-based reconstruction systems that incorporate lidar [9] or depth cameras [18, 17], our key innovation is a rigorous study of how to quantify uncertainty in the resultant reconstruction, which enables improved reconstruction accuracy as well as facilitating downstream tasks such as view selection [19] and navigation within a NeRF map [20]. After training, SiLVR computes the epistemic uncertainty of the NeRF with the Laplace approximation (LA) [21] and Fisher-Information-approximated Hessian matrix. As an efficient alternative to ensemble learning (which requires multiple iterations of training of the NeRF model and is time-consuming), we build upon the formulation of the perturbation field proposed in BayesRays [22] and use the spatial variance of the field as the measure of epistemic uncertainty. The estimated uncertainty is used to filter out reconstruction artefacts which improves the final reconstruction accuracy. This is particularly important when merging NeRF submaps since these submaps often contain artefacts caused by limited observation at the submap boundaries. In addition, we adapt a previously developed lidar SLAM algorithm [23, 24] to bootstrap the SfM component and to enforce an accurate metric scale. For mapping large-scale building complexes or a city block, we adopt a submapping approach and apply a graph partitioning algorithm [25] with visibility information to divide the complete trajectory into submaps. We study how the use of visibility information allows the submaps to be more self-contained and to have fewer artefacts at their boundaries compared to methods that only consider spatial information [9].
@@
  • An uncertainty-aware NeRF reconstruction system fusing lidar and vision that can reconstruct large-scale outdoor environments.
  • Epistemic uncertainty quantification of the multi-modal NeRF pipeline as a principled approach to quantify the contribution of lidar and visual information to a 3D reconstruction, which can be used to improve reconstruction accuracy, especially at submap boundaries. Our method extends the vision-only uncertainty estimation framework proposed in BayesRays [22] to also support lidar depth. This allows us to identify areas with reliable reconstruction (e.g., where there are visual features or abundant lidar measurements) and unreliable areas (e.g., uniformly-textured surfaces which have also not been scanned by lidar).
@@ II Related Work
In this section, we first review classical 3D reconstruction methods based on lidar sensors or cameras, followed by more recent approaches using radiance field representations. We then discuss methods to quantify the uncertainty of a NeRF reconstruction, and techniques that can extend the NeRF methods to large-scale environments.
@@ II-A Classical 3D Reconstruction
Lidars and cameras are the two main modalities used in robotic perception and specifically for 3D reconstruction. For each sensor modality, trajectory estimation is a typical first step in a reconstruction pipeline. In this section, we review classical lidar-based and vision-based pose estimation and 3D reconstruction methods. Then, we review strategies for extending these methods to large-scale scenes.
@@
Lidar is the dominant sensor used for accurate 3D reconstruction of large-scale outdoor environments [27, 28]. It actively transmits laser pulses to measure ranges and, as a result, is accurate even at ranges beyond 100m. With these distance measurements, Lidar odometry typically uses a variant of Iterative Closest Point (ICP), and often integrates high-frequency IMU measurements [29, 30, 31, 23]. Small errors in odometry can accumulate over time, resulting in trajectory drift. This drift can be mitigated when a sensor revisits a previous place and detects loop closure with pose graph optimisation [32] which allows a consistent map to be maintained. After registering all the lidar scans, the (surface) reconstruction problem is then to fuse discrete observations into a map. Example map representations include surfels [33, 34], voxels [35, 36, 37, 38] and wavelets [39]. Despite its advantages, lidar has its own limitations. Lidar sensors are much more expensive than cameras, and their measurements are typically much sparser than camera images. The measurements have inherent noise with ranging errors in the order of centimetres, which makes it difficult to reconstruct small objects accurately in indoor scenes. Finally, lidar data contains no texture or colour, so the final reconstruction is only geometric and cannot be used for applications such as view synthesis, which requires texture.
@@
Alternatively, textured reconstructions can be recovered from camera images alone via SfM. Given the correspondences between images, a SfM system [1] can optimise a set of camera poses, camera intrinsics, and 3D sparse feature points. This can then be used by a MVS system [40] to compute dense depth for each frame and in turn to create a dense 3D point cloud. Compared to lidar, cameras are much more affordable, and also provide texture and colour. However, the performance of visual reconstruction method depends on environmental conditions, and the quality of feature matching. This makes the resultant reconstruction less reliable in scenes that contain repetitive patterns, low-texture surfaces, poor lighting conditions and non-Lambertian materials.
@@
When the scene to be reconstructed is large-scale (e.g., urban districts or multi-room indoor environments), computer memory becomes a limiting factor. Attempting to map a large scene while constraining the output map size might result in a lower-resolution reconstruction lacking detail. A common strategy to extend dense reconstructions to large-scale areas is to divide the scene into submaps [41]. For visual reconstruction with many thousands of images [42], a submapping approach can significantly reduce computation time and memory requirements. One approach used in large-scale MVS is the submap partitioning [43] which groups images into clusters while not degrading the final resultant reconstruction. After partitioning, each submap should be filtered and merged into one unified model. For online lidar mapping systems, the motivation for using submapping techniques is to accommodate loop closure corrections into the already-built map (Occupancy grid or TSDF). These systems construct submaps that are attached to a pose graph [44, 45, 46], and can deform each submap by reoptimising the pose graph upon loop closure.
@@ II-B Radiance Field Representation
Neural Radiance Fields (NeRF) were first proposed in the seminal paper from Mildenhall et al. [3]. The technique uses a multilayer perceptron (MLP) to represent a continuous radiance field, and uses differentiable volume rendering to reconstruct novel views. It minimises the photometric error between the rendered image and the input image, which implicitly achieves multi-view consistency. NeRF and its many variants use frequency encoding [47] to encode spatial coordinates, but these suffer from long training times, typically a few hours per scene. Alternative explicit representations of radiance fields, including dense voxel-grids with trainable per-vertex features [48, 12] and more recently 3D Gaussians [4], are shown to accelerate the training, at the cost of being more memory intensive. Octree or sparse-grid structures [49, 12] can reduce memory usage by pruning grid-features in empty space. Our work is built upon Nerfacto from the open-sourced Nerfstudio project [15]. It incorporates the main features from other NeRF works [12, 50, 51] that have been found to work well with real data.
@@
Neural field representations have been used for lidar-based mapping [55, 56, 57], showing promise in generating more complete and compact reconstructions than traditional methods. While these works also build upon implicit map representations, they do not use visual data to build the map. Our system uses visual information and multi-view geometry constraints. Because of this, it can reconstruct regions outside of the lidar’s FOV.
@@ III-A Radiance Field Formulation
where $\mathbf{c}_{i}$ is the colour of the $i$-th point sample along the ray $\mathbf{p}_{i}$, and its weighting coefficient $w_{i}$ is defined by
@@
where $\delta_{i}$ is the distance between adjacent samples, and $\sigma_{i}$ is the volume density of $\mathbf{p}_{i}$.
@@
[…]
@@ III-B Bayesian Interpretation of the NeRF training
The NeRF reconstruction (i.e., the radiance field function $f$) is deterministic and has no explicit notion of uncertainty. In practice, different parts of the NeRF reconstruction are inherently more uncertain or less reliable. For example, ill-conditioned visual constraints from non-textured areas or insufficient visual parallax can cause the visual reconstruction accuracy to deteriorate. Another example specifically relevant to our work is that a surface is more uncertain if it has only been sparsely scanned by the lidar with limited range and FOV compared to a surface that is densely scanned. Quantifying these uncertainties associated with the NeRF reconstruction allows one to identify the unreliable reconstructions and filter them out accordingly, hence improving reconstruction accuracy. It can also enable downstream tasks, such as view selection for more accurate and complete mapping.
@@
$Z=\int p(\mathcal{D}\mid\theta)p(\theta)\,d\theta$ (6)
Here, the posterior $p(\theta|\mathcal{D})$ is used for the uncertainty estimation described later in Sec. III-C.
III-C Laplace Approximation
Laplace approximation is a technique to estimate the otherwise intractable posterior distribution as a Gaussian function. This allows one to efficiently approximate the posterior distribution, and has been used widely in Bayesian deep learning literature [63]. In this section, we describe the details of the Laplace Approximation, following the presentation by Daxberger et al. [21].